OpenStudy (anonymous):

USE PARTIAL FRACTION to integrate sina/cos^2a+cosa-2

OpenStudy (anonymous):

Set u = cosa so that du = -sinada and then use partial fractions

OpenStudy (anonymous):

\[\int\frac{\sin a}{\cos^2a+\cos a-2}~da\] First substitute \(u=\cos a\), so that \(-du=\sin a~da\). Makes it look easier to work with: \[-\int\frac{du}{u^2+u-2}\]